need this solved in a proportion
Katherine drove 100 miles on 7 gallons of gasoline. How far can she drive on a full tank of 20 gallons?

Answer by nerdybill(7384)

Katherine drove 100 miles on 7 gallons of gasoline. How far can she drive on a full tank of 20 gallons?
.
From:"Katherine drove 100 miles on 7 gallons of gasoline." we get the following proportion:
100/7
.
Because we would like to consider how far 20 gallons will go instead of 7 we have our equation of:
If x = distance traveled on 20 gals of gas
100/7 = x/20
(100/7)20 = x
285.714 miles = x